Problem 3.52 Elina Siljander owns Elinas Stained Glass in Helsinki, Finland. The business produces and sells three differentX Your answer is incorrect. Try again. Calculate Elinas degree of operating leverage for September. (Round to 2 decimal placWhat are the correct answers?
Answer #1

ANSWER

Part 1

Cost function: total cost = €6771 + 28% of total revenue

Raw materials are variable = 4029/16116 = 25%

Labor and rent are fixed

Miscellaneous using regression analysis comes to 289 fixed and 3% variable

SUMMARY OUTPUT

Regression Statistics

Multiple R

0.873015

R Square

0.762156

Adjusted R Square

0.643234

Standard Error

69.9928

Observations

4

ANOVA

df

SS

MS

F

Significance F

Regression

1

31397.01

31397.01

6.408871

0.126985

Residual

2

9797.985

4898.993

Total

3

41,195

Coefficients

Standard Error

t Stat

P-value

Lower 95%

Upper 95%

Lower 95.0%

Upper 95.0%

Intercept

288.5315

162.9787

1.770363

0.218684

-412.709

989.7724

-412.709

989.7724

X Variable 1

0.033137

0.013089

2.531575

0.126985

-0.02318

0.089456

-0.02318

0.089456

Part 2

Revenue = fixed expense + profit / contribution margin ratio = (6771+10000)/(1-0.28) = $23,293

Part 3

Degree of operating leverage = contribution margin / profit = (16116*(1-0.28))/4812 = 2.41

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************